OpenStudy (gahm8684):

How many solutions will this system of equations have? y = 3.5x - 3.5 y = -3.5x + 3.5 No solution Infinite solutions One solution Two solutions

OpenStudy (gahm8684):

I'm giving medals

OpenStudy (anonymous):

one solution..only because it is linear eq and each linear eq has a single solution..okay..??

OpenStudy (gahm8684):

perfect! thanks dude

OpenStudy (anonymous):

rewrite to y-3.5x = -3.5 y+3.5x = 3.5 solve top equation ---> x = y/3.5 +1 Substitute this into bottom equation: y + 3.5(y/3.5 +1) = 3.5 Simplify y + (y + 3.5) = 3.5 2y +3.5 = 3.5 2y = 0 y = 0 Substitute y = 0 in an equation to find x: 0 + 3.5x = 3.5 x = 1 It helps to know why